dear sir, solve for n;(2n+1)C1 here is: out of 2n+1 objects, 1 object selected. ..and not (2n+1)*C1 (2n+1)C1 + (2n+1)C2 + (2n+1)C3 +....+(2n+1)Cn = 63 [answer ; n=3]

dear sir,


solve for n;(2n+1)C1 here is:out of 2n+1 objects, 1 object selected...and not (2n+1)*C1


(2n+1)C1 + (2n+1)C2 + (2n+1)C3 +....+(2n+1)Cn = 63


[answer ; n=3]

Grade:12

3 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear H V,

Ans:- Let us consider the expansion (1+X)^(2n+1)

(1+X)^(2n+1)=(2n+1)C0+(2n+1)C1 X +(2N+1)C2 X²+............+(2N+1)C2N X^2N+(2N+1)C(2N+1)X^2N+1................1

Now we know that NCr=NC(N-r)

(2N+1)C1=(2n+1)C2N

(2N+1)C2=(2N+1)C(2N-1)       

Again ,

(2N+1)CN=(2N+1)C(N+1) etc

Hence from these relations we get,(Putting X=1 in eq 1)

2^(2N+1)=(2N+1)C0+2K+(2N+1)C(2N+1)

where K=(2N+1)C1+(2N+1)C2+..............(2N+1)C(2N+1)=63(given)

2^(2N+1)=1+1+126=128

4^N=64=4³

Hence N=3(Ans) 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best H V !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur

h v
13 years ago

can anyone give the answer please?

sunil
6 years ago
formula:
(2n+1)C0 +(2n+1)C1+ (2n+1)C2 + (2n+1)C3 +....+(2n+1)Cn=1/2{(2n+1)C0 +(2n+1)C1+ (2n+1)C2 + (2n+1)C3 +....+(2n+1)C2n+1} 
again (2n+1)C0 + (2n+1)C2 + (2n+1)C3 +....+(2n+1)C2n+1= 2^(2n+1)
so (2n+1)C0 +(2n+1)C1+ (2n+1)C2 + (2n+1)C3 +....+(2n+1)Cn=1/2(2^(2n+1))....eq 1
given:
(2n+1)C1 + (2n+1)C2 + (2n+1)C3 +....+(2n+1)Cn = 63
put this in eq1
so (2n+1)C0 +63=1/2(2^(2n+1))
64=1/2(2^(2n+1)) as (2n+1)C0=1
solve this n comes out to be 3